# Evolver: Chain-of-Evolution Prompting to Boost Large Multimodal Models for Hateful Meme Detection

### v1

@@ Abstract.
Recent advances show that two-stream approaches have achieved outstanding performance in hateful meme detection. However, hateful memes constantly evolve as new memes emerge by fusing progressive cultural ideas, making existing methods obsolete or ineffective. In this work, we explore the potential of Large Multimodal Models (LMMs) for hateful meme detection. To this end, we propose Evolver, which incorporates LMMs via Chain-of-Evolution (CoE) Prompting, by integrating the evolution attribute and in-context information of memes. Specifically, Evolver simulates the evolving and expressing process of memes and reasons through LMMs in a step-by-step manner. First, an evolutionary pair mining module retrieves the top-$K$ most similar memes in the external curated meme set with the input meme. Second, an evolutionary information extractor is designed to summarize the semantic regularities between the paired memes for prompting. Finally, a contextual relevance amplifier enhances the in-context hatefulness information to boost the search for evolutionary processes. Extensive experiments on public FHM, MAMI, and HarM datasets show that CoE prompting can be incorporated into existing LMMs to improve their performance. More encouragingly, it can serve as an interpretive tool to promote the understanding of the evolution of social memes. [^1] The source code will be made publicly available. Disclaimer. This paper contains offensive content that may be disturbing to some readers.
@@
large multimodal models, hateful meme detection, prompting
price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Scene understanding
1. Introduction
@@
Figure 1. The illustration of (a) the evolution of memes and comparison between (b) conventional two-stream methods, and (c) our Evolver method. Memes evolve by fusing new cultural concepts. The meme of Trump is influenced by the meme of a sad frog in an image and text symbol, which creates a new hateful meme. Conventional hateful meme detection methods use trainable two-stream encoders and fusion for meme classification, with poor interpretability. In contrast, our Evolver captures the evolution and context of memes, utilizing them as prompts for large multimodal models to obtain a comprehensive understanding of memes.
@@
Hateful meme detection (Lippe et al., 2020; Kiela et al., 2020; Cao et al., 2020) is a crucial task in the field of multimodal research, aiming to identify content that combines text and images to propagate hate speech or offensive messages. Memes, as a widespread cultural phenomenon, proliferate on the Internet, blending images and texts to convey sophisticated meanings. Furthermore, the fusion of visual and textual elements in memes complicates the interpretation of their semantics and the identification of hateful undertones. The combination of text and images presents significant challenges in hateful meme detection, especially for detecting and moderating hateful content (Levine, 2013).
@@
With the advances in image-text pre-training, efforts to leverage and fine-tune CLIP (Radford et al., 2021) for hateful meme detection have demonstrated notable success in achieving high accuracy (Kumar and Nandakumar, 2022; Arya et al., 2024; Hee et al., 2022). However, as shown in Figure 1(a), memes evolve similarly to biological organisms (Dawkins, 2016b, a), undergoing variation, mutation, and inheritance, which complicates the detection of hateful memes as they continuously evolve, merging with other cultural elements to form new expressions. Traditional methods for hateful meme detection, as illustrated in Figure 1(b), suffer from limitations in interpretability and adaptability. These shortcomings hinder the understanding of memes’ evolving nature and their contextual nuances, leading to overfitting on training datasets and diminished effectiveness in the dynamic and evolving real-world meme landscape.
@@
To overcome this dilemma, we introduce Evolver, a novel approach that uses Large Multimodal Models (LMMs) (Li et al., 2023a; Lyu et al., 2023; Yang et al., 2023) for hateful meme detection. By developing a benchmark specifically tailored for hateful meme detection based on LMMs, we aim to address the aforementioned limitations of traditional detection methods. Our approach not only enhances detection capabilities but also provides a more explainable and adaptable framework suitable for the evolving landscape of Internet memes. As shown in Figure 1(c), Evolver is an innovative approach to advance the state-of-the-art in zero-shot hateful meme detection significantly. Extensive experiments have shown the effectiveness of our method across three widely recognized hateful meme detection datasets, demonstrating a superior ability to identify and interpret hateful memes.
@@
  • Extensive experiments on commonly used zero-shot hateful meme detection benchmarks with superior performance validate the efficacy and generalization of our method.
2. Related Work
Hateful Meme Detection.    Most works of hateful meme detection are based on pretrained vision-language models and fine-tune them on the hateful meme detection data. For example, Kiela et al. (2020) introduced the Hateful Memes Challenge and proposed a multimodal model that combines visual and textual features using a transformer-based architecture. Hate-CLIPper proposed a multimodal contrastive learning approach to improve the representation learning of memes by aligning the visual and textual modalities. These works demonstrate the effectiveness of using pretrained vision-language models. Differently, we discuss the potential of LMMs for hateful meme detection and introduce a knowledge-enhance LMM, Evolver.
Large Multimodal Models.    With the release of the GPT-4V (OpenAI, 2023), many researchers focus on the development and application of LMMs in processing and understanding multimodal data. Discuss groundbreaking models like GPT-4V and LLAVA-1.5 (Liu et al., 2023a), highlighting their capabilities in interpreting complex datasets that include a combination of text and images. Address the challenges these models face, such as sensitivity to noise and typographical errors, and the ongoing efforts to improve their accuracy and robustness in real-world applications. In this work, we adopt the LMMs to hateful meme detection, where set an LMM-based zero-shot hateful meme detection benchmark. Furthermore, we show that our Evolver further boosts the LMM’s performance.
3. Approach
In this section, we first briefly describe the hateful meme detection task and LMM-based approaches. Next, we introduce the proposed Evolver.
3.1. Background
3.1.1. Problem Definition
@@
3.1.2. Large Multimodal Models
Vision Encoder    serves as a translator to help language models understand visual content. It leverages frozen pretrained vision models such as CLIP (Radford et al., 2021) and ViT (Dosovitskiy et al., 2020) to encode visual content so that the language model can understand visual content:
(2) $h_{v}=\bm{W}\cdot{\text{Enc}_{vis}}(X_{v}),$
@@ 3.1.2. Large Multimodal Models
Large Language Model Decoder    generates a sentence given tokens (Touvron et al., 2023). The process of generation can be represented as:
(3) $h_{t}=\text{Tokenize}(X_{t}),$
(4) $p(w)=\prod^{n}_{i=1}p(w_{i}|w_{<i},h_{t},h_{v}),$
@@
3.2. Evolver: Chain-of-Evolution Prompting
@@
Evolutionary Pair Mining    It is unusual to find memes that have an evolutionary relationship with existing memes. Motivated by Qu et al. (2023), the evolution of a meme is defined as new memes that emerge by fusing other memes or cultural ideas. Therefore, the evolution of memes and old memes share similar textual and visual semantic regularities. We leverage this property to identify these old memes given the evolved meme. To this end, we can leverage the evolution of hateful memes to enhance LMMs’ ability to understand hateful memes. In the implementation, we first generate the textual and visual embeddings from an external meme pool and target memes with CLIP and then fuse the textual and visual embeddings with a fixed ratio. Ideally, meme evolution has temporal ability, which means that the memes change with time. However, it is difficult to accurately locate the upstream and downstream of meme evolution with existing technology. Therefore, we use multiple evolutionary memes to find the common characteristics of this evolution. Finally, given a target image-text pair, we retrieve the top-K similar memes using cosine similarity:
(5) $\text{memes}=\{A_{i}|\text{cos}(\mathbf{A},\mathbf{B})_{i}\in\text{Top}_{K}(\text{cos}(\mathbf{A},\mathbf{B}))\},$
@@
(6) $\text{Info}=\text{LMM}([\text{memes}_{top_{K}},X_{extract}]),$
where Info stands for our evolutionary information and LMM indicates the large multimodal model. $\text{memes}_{Top_{K}}$ are the $Top_{K}$ memes retrieved in the previous step, and $X_{extract}$ is the instruction to guide the LMM to extract information. We present the detailed instruction of $X_{extract}$ as follows:
@@
Input: [image 0 : ¡image0¿, caption 0 : texts[0], image 1 : ¡image1¿, caption 1 : texts[1], image 2 : ¡image2¿, caption 2 : texts[2], , image 3 : ¡image3¿, caption 3 : texts[3], image 4 : ¡image4¿, caption 4 : texts[4]’]
@@
Table 1. The example of the instruction for extracting the evolutionary information.
@@
(7) $\hat{y}=\text{LMM}([\text{memes}_{target},X_{detect},\text{Info},\text{Amplifier}]),$
where Info stands for the information extracted previously, $\hat{y}$ is the final prediction, $\text{memes}_{target}$ is the memes we want to detect, and $X_{detect}$ is the instruction to ask the large multimodal model to detect hateful memes. Amplifier refers to contextual relevance amplifier. The example of the amplifier is the same as the blue part in Table 1.
@@
Gemini-Pro-V (Team et al., 2023) - 66.0 65.7 74.5 74.5 71.3 76.2
GPT-4V (OpenAI, 2023) - 70.5 70.3 - - - -
@@
Openflamingo (Awadalla et al., 2023) 7B 57.0 56.4 56.8 56.8 51.7 55.8
LLaVA-1.5 (Liu et al., 2023a) 13B 61.8 61.4 57.4 57.4 55.0 54.5
MMICL (Zhao et al., 2023) 11B 59.9 60.4 67.3 67.3 52.1 63.8
MiniGPT-v2 (Zhu et al., 2023) 7B 58.8 59.1 62.3 62.3 57.1 60.3
BLIP-2 (Li et al., 2023b) 11B 56.4 55.8 59.4 59.4 56.8 60.6
InstructBLIP (Dai et al., 2023) 13B 59.6 60.1 64.1 64.1 55.7 60.1
@@
Evolver (Ours) 13B 60.3 60.4 59.9 59.9 59.3 57.3
Table 2. Comparison among different LMMs on zero-shot hateful meme detection benchmarks. 11B and 13B Evolver are based on MMICL and LLaVA-1.5, respectively. The API-based models have ethical considerations, we can not directly apply our CoE prompt above them. Best results in the open-source setting are highlighted in bold.
@@
Table 3. Ablation study of the three components of our method on FHM dataset.
Figure 2. Example results of the Evolver (Ours) and the baseline model (MMICL). More examples refer to the Appendix 5.
4. Experiment
We first describe three public datasets as well as the widely-used metrics and settings. Next, we introduce the baseline models including API-based (e.g., GPT-4V) and open-source LMMs (e.g., LLaVA) in Appendix LABEL:sec:baseline. Finally, we present the evaluation and ablation study of our CoE prompting.
@@ 4. Experiment
GPT-4V    (OpenAI, 2023) first creates a GPT-based LMM on the massive volume of datasets on the internet. They trained the language decoder, then aligned it with the vision encoder.
Gemini-Pro-V    (Team et al., 2023) curates a large multimodal model with various image-text data on the web and then trains the model with image-text data from scratch.
Openflamingo    (Awadalla et al., 2023) builds a LMM using CLIP as vision encoder and MPT/RedPajama as language decoder. The model is trained on open-source image-text datasets.
LLaVA1.5    (Liu et al., 2023a) improves the performance of the model they propose previously (Liu et al., 2023b) where CLIP is used as vision encoder and Vicuna as language decoder, with MLP projection to bridge the gap between vision encoder and language decoder. The model is pre-trained on the curated collected dataset. Furthermore, the model is instructionally fine-tuned with academic-task-oriented VQA data.
MiniGPTv2     (Zhu et al., 2023) connects the vision and language space with Q-former and an MLP projection. The model is pre-trained with a large collection of aligned image-text pairs and then instruction fine-tuned on curated high-quality image-text pairs.
BLIP-2     (Li et al., 2023b) propose a model architecture using only Q-former to connect the vision encoder and language decoder to perform a series of multi-modal tasks. The pre-training of model has two stages. The initial stage initiates the learning of vision-language representations using a frozen vision encoder. Subsequently, the second stage advances the generative learning of vision-to-language from a stationary language model.
InstructBLIP     (Dai et al., 2023) further performs instruction fine-tuning based on BLIP-2 various converted multi-model datasets.
MMICL    (Zhao et al., 2023) propose a framework using Q-former and an MLP layer to connect a vision encoder and language decoder.
4.1. More Experiment Details
Settings    We implement our Evolver based on the MMICL and LLaVA-1.5. For MMICL, we set the minimum length for the generation to 50, the maximum length for the generation to 80 during the Evolution Info Extractor, and the temperature to 0.2. During the final prediction, we set the minimum length of generation to 1, and the maximum length for generation to 50. For LLaVA-1.5, we set temperature to 0.2 and maximum generated tokens to 1024 for both stages.
Effect of the Ratio of Textual and Visual fusion.    The evolution of memes should consider both textual and visual information in a mixed form. We generate the textual and visual embeddings from the carefully curated meme pool and target memes with CLIP and then fuse the textual and visual embeddings with a 4:1 ratio in practice.
Effect of the External Curated Meme Pool.    Generally, the external curated meme pool should have two characteristics (1) do not overlap with the test set. (2) have enough evolutionary information. To simplicity, we use the training set as the carefully curated meme pools, where the memes follow the same definition of hatefulness/harm/misogyny.
Dataset Details.    For more clarity, we show the detailed hateful meme detection dataset analysis in Table 4. Here, we give a detailed description of each dataset.

### v2

@@ Abstract
Hateful memes continuously evolve as new ones emerge by blending progressive cultural ideas, rendering existing methods that rely on extensive training obsolete or ineffective. In this work, we propose Evolver, which incorporates Large Multimodal Models (LMMs) via Chain-of-Evolution (CoE) Prompting, by integrating the evolution attribute and in-context information of memes. Specifically, Evolver simulates the evolving and expressing process of memes and reasons through LMMs in a step-by-step manner using an evolutionary pair mining module, an evolutionary information extractor, and a contextual relevance amplifier. Extensive experiments on public FHM, MAMI, and HarM datasets show that CoE prompting can be incorporated into existing LMMs to improve their performance. More encouragingly, it can serve as an interpretive tool to promote the understanding of meme evolution.
Disclaimer. This paper contains offensive content that may be disturbing to some readers.
@@
Evolver: Chain-of-Evolution Prompting to Boost Large Multimodal Models for Hateful Meme Detection
Jinfa Huang1, Jinsheng Pan1thanks: Equal contribution., Zhongwei Wan2, Hanjia Lyu1, Jiebo Luo1 1University of Rochester  2The Ohio State University {jhuang90, jpan24, hlyu5}@ur.rochester.edu, wan.512@osu.edu,  jluo@cs.rochester.edu https://github.com/inFaaa/Evolver
1 Introduction
@@
Figure 1: The illustration of (a) the evolution of memes and comparison between (b) conventional two-stream methods, and (c) our Evolver method. Memes evolve by fusing new cultural concepts. The meme of Trump is influenced by the meme of a sad frog in an image and text symbol, which creates a new hateful meme. Conventional hateful meme detection methods use trainable two-stream encoders and fusion for meme classification, with poor interpretability. In contrast, our Evolver captures the evolution and context of memes, utilizing them as prompts for large multimodal models to obtain a comprehensive understanding of memes.
@@
Hateful meme detection Lippe et al. (2020); Kiela et al. (2020); Cao et al. (2020) is a crucial task in the field of multimodal research, aiming to identify content that combines text and images to propagate hate speech or offensive messages. Memes, as a widespread cultural phenomenon, proliferate on the Internet, blending images and texts to convey sophisticated meanings. Furthermore, the fusion of visual and textual elements in memes complicates the interpretation of their semantics and the identification of hateful undertones. The combination of text and images presents significant challenges in hateful meme detection, especially for detecting and moderating hateful content Levine (2013).
@@
With the advances in image-text pre-training, efforts to leverage and fine-tune CLIP Radford et al. (2021) for hateful meme detection have demonstrated notable success in achieving high accuracy Kumar and Nandakumar (2022); Arya et al. (2024); Hee et al. (2022). However, as shown in Figure 1(a), the extensive evolution of memes fusing together complicates the detection of hateful memes as they continuously evolve, bring new cultural elements to form new expressions. Traditional methods for hateful meme detection, as illustrated in Figure 1(b), suffer from limitations in interpretability and adaptability. These shortcomings hinder the understanding of memes’ evolving nature and their contextual nuances, leading to overfitting on training sets and diminished effectiveness in the dynamic and evolving meme landscape.
@@
To overcome this dilemma and inspired by  Dawkins (2016b, a) which shows that similar biological groups share some common traits, we introduce Evolver, a novel approach that uses Large Multimodal Models (LMMs) Li et al. (2023a); Lyu et al. (2023); Yang et al. (2023) for hateful meme detection. By developing a benchmark specifically tailored for hateful meme detection based on LMMs, we aim to address the aforementioned limitations of traditional detection methods. Our approach not only enhances detection capabilities but also provides a more explainable and adaptable framework suitable for the evolving landscape of Internet memes. Evolver is an innovative approach in zero-shot hateful meme detection significantly. Extensive experiments have shown the effectiveness of our method across three widely recognized hateful meme detection datasets, demonstrating a superior ability to identify and interpret hateful memes.
@@
  • Extensive experiments on commonly used zero-shot hateful meme detection datasets with superior performance validate the efficacy and generalization of our method.
2 Approach
2.1 Problem Definition
@@
2.2 Large Multimodal Models
Vision Encoder    serves as a translator to help language models understand visual content. It leverages frozen pretrained vision models such as CLIP Radford et al. (2021) and ViT Dosovitskiy et al. (2020) to encode visual content so that the language model can understand visual content:
$h_{v}=\bm{W}\cdot{\text{Enc}_{vis}}(X_{v})$ (2)
@@ 2.2 Large Multimodal Models
Large Language Model Decoder    generates a sentence given tokens Touvron et al. (2023). The process of generation can be represented as:
$h_{t}=\text{Tokenize}(X_{t})$ (3)
$p(w)=\prod^{n}_{i=1}p(w_{i}|w_{<i},h_{t},h_{v})$ (4)
@@
2.3 Evolver: Chain-of-Evolution Prompting
@@
Evolutionary Pair Mining    It is unusual to find memes that have an evolutionary relationship with existing memes. Motivated by Qu et al. (2023), the evolution of a meme is defined as new memes that emerge by fusing other memes or cultural ideas. Therefore, the evolution of memes and old memes share similar textual and visual semantic regularities. We leverage this property to identify these old memes given the evolved meme. To this end, we can leverage the evolution of hateful memes to enhance LMMs’ ability to understand hateful memes. In the implementation, we first generate the textual and visual embeddings from an external meme pool and target memes with CLIP. Specifically, the external meme pool should have two characteristics (1) do not overlap with the test set. (2) have enough evolutionary information. For simplicity, we use the training set as the carefully curated meme pools rather than any other dataset, where the memes follow the same definition of hatefulness/harm/misogyny. Then, we fuse the textual and visual embeddings with a fixed ratio. Ideally, meme evolution has temporal ability, which means that the memes change with time. However, it is difficult to accurately locate the upstream and downstream of meme evolution with existing technology. Therefore, we use multiple evolutionary memes to find the common characteristics of this evolution. Finally, given a target image-text pair, we retrieve the top-K similar memes using cosine similarity:
$\text{memes}=\{A_{i}|\text{cos}(\mathbf{A},\mathbf{B})_{i}\in\text{Top}_{K}(\text{cos}(\mathbf{A},\mathbf{B}))\}$ (5)
@@
$\text{Info}=\text{LMM}([\text{memes}_{top_{K}},X_{extract}])$ (6)
where Info stands for our evolutionary information and LMM indicates the large multimodal model. $\text{memes}_{Top_{K}}$ are the $Top_{K}$ memes retrieved in the previous step, and $X_{extract}$ is the instruction to guide the LMM to extract information. We present the detailed instruction of $X_{extract}$ as shown in Table 1. More definitions of hateful memes from different datases are shown in Appendix D.
@@
Input: [image 0 : <image0>, caption 0 : texts[0], image 1 : <image1>, caption 1 : texts[1], image 2 : <image2>, caption 2 : texts[2], , image 3 : <image3>, caption 3 : texts[3], image 4 : <image4>, caption 4 : texts[4]’]
@@
Table 1: The example of the instruction for extracting the evolutionary information.
@@
$\hat{y}=\text{LMM}([\text{memes}_{T},X_{D},\text{Info},\text{Amp}])$ (7)
where Info stands for the information extracted previously, $\hat{y}$ is the final prediction, $\text{memes}_{T}$ is the memes we want to detect, and $X_{D}$ is the instruction to ask the large multimodal model to detect hateful memes. Amp refers to contextual relevance amplifier. The example of the amplifier is the same as the blue part in Table 1.
The Principles of Prompts    Our prompt strategy is simple yet effective, based on two key principles: (1) include the hateful meme definition and (2) limit the prompt to 30 words, addressing LMMs’ challenges with long-text comprehension. We directly applied definitions from the FHM, MAMI, and HarM datasets, summarized to meet this length requirement with the whole prompt using GPT-4. To this end, we do not use any specialized prompt design, highlighting the robustness of our method.
@@
Gemini-Pro-V Team et al. (2023) - 66.0 65.7 74.5 74.5 71.3 76.2
GPT-4V OpenAI (2023) - 70.5 70.3 - - - -
@@
Openflamingo Awadalla et al. (2023) 7B 57.0 56.4 56.8 56.8 51.7 55.8
LLaVA-1.5 Liu et al. (2023a) 13B 61.8 61.4 57.4 57.4 55.0 54.5
MMICL Zhao et al. (2023) 11B 59.9 60.4 67.3 67.3 52.1 63.8
MiniGPT-v2 Zhu et al. (2023) 7B 58.8 59.1 62.3 62.3 57.1 60.3
BLIP-2 Li et al. (2023b) 11B 56.4 55.8 59.4 59.4 56.8 60.6
InstructBLIP Dai et al. (2023) 13B 59.6 60.1 64.1 64.1 55.7 60.1
@@
Evolver (Ours) 13B 62.3 62.5 59.9 59.9 59.3 57.3
Table 2: Comparison among different LMMs on zero-shot hateful meme detection benchmarks. Evolver and Evolver are MMICL and LLaVA 1.5 with CoE prompting. The API-based models have ethical considerations, we can not directly apply our CoE prompt above them. Best results in the open-source setting are highlighted in bold.
@@
Table 3: Ablation study of the three components of our method on FHM dataset.
Figure 2: Example results of the Evolver (Ours) and the baseline model (MMICL). For more examples refer to the Appendix.
3 Experiment
3.1 Experimental Setup
Datasets    We evaluate Evolver on three widely-used public datasets, namely, Facebook Hateful Meme dataset (FHM) Kiela et al. (2020), Harmful Meme dataset (HarM) Pramanick et al. (2021), and Multimedia Automatic Misogyny Identification (MAMI) Fersini et al. (2022).
Implementation    We implement our Evolver based on the MMICL and LLaVA-1.5. For MMICL, we set the minimum length for the generation to 50, the maximum length for the generation to 80 during the Evolution Info Extractor, and the temperature to 0.2. During the final prediction, we set the minimum length of generation to 1, and the maximum length for generation to 50. For LLaVA-1.5, we set temperature to 0.2 and maximum generated tokens to 1024 for both stages. The embedding size of textual and visual embeddings is $N$$\times$ 768, and $N$ is the number of memes. We fuse textual and visual embeddings with a fixed ratio of 4:1 by element-wise add in practice.
Baselines    The detailed description of the baseline models including API-based (e.g., GPT-4V) and open-source LMMs (e.g., LLaVA) can be found in Appendix A.
LMM Backbones    We implement our Evolver based on MMICL Zhao et al. (2023) and LLaVA-1.5 Liu et al. (2023a). Please refer to the Appendix B for more details.
3.2 Main Result
Table 2 highlights the effectiveness of our Chain-of-Evolution Prompting strategy across three datasets. We compare zero-shot results among closed-source LMMs, open-source LMMs, and open-source LMMs with the CoE Prompting strategy. First, we observe that with more evolutionary context, LMMs exhibit a greater ability to recognize hateful content. Notably, both MMICL and LLaVA-1.5 with CoE Prompting outperform their zero-shot baselines, with LLaVA-1.5 achieving a 1.1% improvement and MMICL a 3.2% improvement in accuracy on the FHM dataset. Additionally, within open-source LMMs, model size does not necessarily lead to a better understanding of hateful memes. For instance, on MAMI, MiniGPT-v2 (7B) outperforms LLaVA-1.5 (13B). Furthermore, despite some closed-source LMMs being unavailable for evaluation on certain datasets due to model safety updates, these closed-source models still outperform open-source LMMs in zero-shot results, suggesting a significant gap for open-source models to close. The typical existing training-based models are fully supervised settings, mainly based on CLIP Radford et al. (2021) and BERT Devlin (2018), which can not provide interpretability. However, our setting is a zero-shot setting, mainly including the LMM-based methods, which can explicitly reflect the evolution of memes and training-free. The superior performance of Gemini-Pro-V and GPT-4V is due to their larger parameter count and extensive training data. Furthermore, without security restrictions, our method could be seamlessly integrated into these models.
3.3 Ablation Study
3.4 Qualitative Analysis
We visualize the paired evolution memes retrieved by the evolutionary pair mining modules in Figure 2, showing how the evolution information influences the final prediction. The origin model (MMICL) without considering the meme’s evolution, predicts non-hateful because it is not easy to detect the boy’s disability or genetic conditions. After obtaining the evolution information like “discriminates against individuals with disabilities or genetic conditions.", our Evolver rectify the prediction, which supports the rationale of our method.
3.5 Impact of Number of Evolutionary Memes
Figure 3: Effect of the number of evolutionary memes.
4 Conclusion
We present Evolver to seamlessly boost LMMs for hateful meme detection via Chain-of-Evolution prompting. By integrating evolution of memes, our method can adapt to unseen memes. Experimental results demonstrate the effectiveness of Evolver.
Impact Statements
Ethics Statement.    While our objective is to mitigate the spread and impact of hate speech online, we recognize the potential for misuse or unintended consequences of this technology. Our framework is designed exclusively for detecting hateful memes, with its use strictly limited to academic or approved research environments. It is not intended for content generation. Insights from this research can support public awareness campaigns, encouraging informed digital citizenship and empowering users to actively participate in recognizing and reporting harmful content. We are committed to ensuring that our technology is used ethically and responsibly.
Appendix A Baseline Details
@@
GPT-4V    OpenAI (2023) first creates a GPT-based LMM on the massive volume of datasets on the internet. They trained the language decoder, then aligned it with the vision encoder.
Gemini-Pro-V    Team et al. (2023) curates a large multimodal model with various image-text data on the web and then trains the model with image-text data from scratch.
Openflamingo    Awadalla et al. (2023) builds a LMM using CLIP as vision encoder and MPT/RedPajama as a language decoder. The model is trained on open-source image-text datasets.
LLaVA 1.5    Liu et al. (2023a) improves the performance of the model they propose previously Liu et al. (2023b) where CLIP is used as vision encoder and Vicuna as language decoder, with MLP projection to bridge the gap between vision encoder and language decoder.
MiniGPTv2     Zhu et al. (2023) connects the vision and language space with Q-former and an MLP projection. The model is pre-trained with a large collection of aligned image-text pairs and then instruction fine-tuned on curated high-quality image-text pairs.
BLIP-2     Li et al. (2023b) propose a model architecture using only Q-former to connect the vision encoder and language decoder to perform a series of multi-modal tasks. The pre-training of model has two stages.
InstructBLIP     Dai et al. (2023) further performs instruction fine-tuning based on BLIP-2 various converted multi-model datasets.
MMICL    Zhao et al. (2023) propose a framework using Q-former and an MLP layer to connect a vision encoder and language decoder.
Appendix B Data Details
For more clarity, we show the detailed hateful meme detection dataset analysis in Table 4. Here, we give a detailed description of each dataset.